• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 00:48:28 on October 28, 2008

    Entry number 245019

    keyword=target activities Monday

    This morning, Jin/Chiranjib checked the transverse line laser/optics. Polarization of both S and P lines were good (up 90%) before the last two mirrors. Also checked the powers of the S and P lines. Both seems reasonable. It's difficult to measure the polarization after the last two mirrors. Ye took out the broken right down-stream pick-up coil and found two wires broken. He repaired the coil(soldered the two wires). Joe installed the reference cell (later tested). Adjusted cooling jets. In the evening, Ye/Ye' re-installed the fixed pickup coil and made adjustment and measurement. Jin/Chiranjib checked the vertical lines (power, polarization and alignment). Polarizations before the last two mirrors are fine. After the last two mirrors, the P line is fine while the S line has significant degradation in polarization. Adjusting the quarter-wave plate does not improve much. We checked the power and alignment and they are reasonable. (Except the N6 line has a significant loss (known problem). We put full power (25 W on each of the first lines N7, N8, N9 and N10). Started optical pumping on the vertical line. Turned on cooling jets (36). Ye worked on EPR fiber and got good amount EPR light. Some details will be filled in by these people (see target log).